# Humans can learn to detect AI-generated texts, or at least learn when they can’t

### v2

@@ Abstract
We used GPT-4o to generate several hundred texts across various genres and text types comparable to Koditex, a multi-register corpus of human-written texts. We then presented randomized text pairs to 255 Czech native speakers who identified which text was human-written and which was AI-generated. Participants were randomly assigned to two conditions: one receiving immediate feedback after each trial, the other receiving no feedback until experiment completion. We recorded accuracy in identification, confidence levels, response times, and judgments about text readability along with demographic data and participants’ engagement with AI technologies prior to the experiment.
@@ 1 Introduction
Participants (n=255) were randomly assigned to two experimental conditions. One group received immediate feedback after each trial, indicating correctness, whereas the second group received no immediate feedback and learned their overall results only upon completing the experiment.
@@ 2.4 Participants
The main sample was primarily recruited from the participant pool managed by the LABELS psycholinguistic laboratory. The pool consists of volunteers interested in participating in linguistic experiments and students who were offered university course credits for their participation. The dataset also includes several participants from the general public who responded to Facebook posts by the experimenters, accounting for approximately 22 participants.
The complete sample comprised 291 participants. However, 33 participants were excluded for failing to answer all attention-check questions correctly (the experiment was demanding, and some students participating for course credits showed insufficient attention). An additional 3 participants were excluded for not being native Czech speakers.
The final sample consisted of 255 participants (female: 181, male: 70, other or preferred not to say: 4), the mean age was 24.08, with participants ranging from 17 to 80 years old. The majority of participants had a bachelor’s (99) or master’s (135) degree as their highest level of education, 14 participants were doctoral graduates, and only a few had completed only highschool (4), vocational (2) or other (1) type of education. The majority of participants had an educational background in social sciences and humanities (107), substantially represented were also students and graduates of natural sciences (75). 46 participants specifically indicated that they have philological education, 11 participants studied or graduated in computer science, and 16 participants selected the option ‘Other’.
@@ 3.1 Correctness of the Answer
We observed a substantial and statistically significant difference between feedback and no feedback groups. The no feedback group showed considerably lower performance, with an average correctness rate of 55.5% (95% CI: 52.0%–58.4%), while the feedback group achieved a higher average success rate of 65.1% (95% CI: 62.2%–67.7%). These findings confirm hypotheses 1 and 2: participants were able to distinguish AI-generated texts from human-written ones at above-chance levels, though no feedback group performed only slightly above chance.
@@ 3.2 Mixed effects logistic regression
-2.25in0in     Conf. Interval Conf. Interval Variable     Estimate SE Lower Upper Exp(B) Lower Upper z p (Intercept)     1.0734 0.3494 0.38859 1.75812 2.925 1.475 5.802 3.072 0.002 Confidence     -0.1387 0.0476 -0.23209 -0.04537 0.87 0.793 0.956 -2.912 0.004 More readable is human     2.0513 0.2193 1.6216 2.48105 7.778 5.061 11.954 9.356 <0.001 Trial order     -0.0124 0.0112 -0.03427 0.00956 0.988 0.966 1.01 -1.105 0.269 Reaction time log     -0.085 0.082 -0.24565 0.07564 0.919 0.782 1.079 -1.037 0.300 Dim. Euclidean distance     0.0988 0.0503 0.00023 0.19729 1.104 1.0 1.218 1.964 0.049 Dim1 difference     -0.2985 0.0849 -0.46501 -0.13208 0.742 0.628 0.876 -3.515 <0.001 Dim2 difference     -0.353 0.135 -0.61764 -0.08829 0.703 0.539 0.915 -2.614 0.009 Dim3 difference     0.0977 0.0415 0.01635 0.17914 1.103 1.016 1.196 2.354 0.019 Dim4 difference     0.1191 0.0495 0.02202 0.21621 1.127 1.022 1.241 2.405 0.016 Dim5 difference     -0.0925 0.0543 -0.1989 0.01386 0.912 0.82 1.014 -1.705 0.088 Dim6 difference     0.0109 0.0345 -0.05665 0.07853 1.011 0.945 1.082 0.317 0.751 Dim7 difference     0.0358 0.0453 -0.05302 0.12466 1.036 0.948 1.133 0.79 0.429 Dim8 difference     -0.0411 0.0343 -0.10839 0.02614 0.96 0.897 1.026 -1.198 0.231 Age     0.0184 0.0138 -0.00865 0.04548 1.019 0.991 1.047 1.334 0.182 Education     0.0777 0.125 -0.16728 0.32262 1.081 0.846 1.381 0.621 0.534 Gender male - female     -0.0203 0.1838 -0.38055 0.34005 0.98 0.683 1.405 -0.11 0.912 Gender other - female     -0.4545 0.5167 -1.4673 0.5583 0.635 0.231 1.748 -0.88 0.379 Reading disorder     0.713 0.6935 -0.64615 2.0722 2.04 0.524 7.942 1.028 0.304 Study field: human - cs     -0.3623 0.3684 -1.08427 0.35974 0.696 0.338 1.433 -0.983 0.325 Study field: nature - cs     -0.2746 0.3896 -1.03812 0.48898 0.76 0.354 1.631 -0.705 0.481 Study field: other - cs     -0.3041 0.4565 -1.19894 0.59068 0.738 0.302 1.805 -0.666 0.505 Study field: philology - cs     -0.4767 0.4133 -1.28683 0.33337 0.621 0.276 1.396 -1.153 0.249 Frequency of AI usage     0.0177 0.0619 -0.10358 0.13901 1.018 0.902 1.149 0.286 0.775 AI attitudes: improvement     0.0842 0.0733 -0.05946 0.22783 1.088 0.942 1.256 1.149 0.251 AI attitudes: risk     -0.0261 0.0578 -0.13944 0.08721 0.974 0.87 1.091 -0.452 0.652 AI attitudes: quality     -0.012 0.0666 -0.14242 0.11845 0.988 0.867 1.126 -0.18 0.857 AI attitudes: creativity     -0.0113 0.0651 -0.13889 0.11627 0.989 0.87 1.123 -0.174 0.862 AI attitudes: understanding     0.0741 0.0627 -0.04882 0.19712 1.077 0.952 1.218 1.182 0.237 AI attitudes: grammar     0.2308 0.0644 0.10451 0.35703 1.26 1.11 1.429 3.582 <0.001 AI attitudes: avoidance     -0.0743 0.0575 -0.18704 0.0385 0.928 0.829 1.039 -1.291 0.197 AI attitudes: dumbing     0.0533 0.0519 -0.04842 0.15492 1.055 0.953 1.168 1.027 0.305 AI attitudes: fascination     -0.0141 0.0811 -0.17312 0.14494 0.986 0.841 1.156 -0.174 0.862
@@
Table notes: Estimates and exponentiated coefficients (Exp(B)) are reported with 95% confidence intervals. Bold p-values indicate statistical significance (p ¡ 0.05).
@@ 3.2.1 Trial order
This variable is important for our first and second research question. If we are asking whether the group that received feedback differs from the one that did not, we assume that those with feedback were able to learn. Their abilities should therefore improve over time. The model confirmed that they do. The variable Trial order reached level of significance p < 0.001 in feedback group (CI 0.027–0.079). The log odds ratio estimate was at 0.053, that means, the later the trial, the higher probability that the participant answers correctly. In the no feedback group, on the other hand, the trial order did not reach the level of significance (p=0.27, estimate -0.012).
FIG 4 shows that the correctness rises rather chaotically, which is due to the high variability of the data (half of 155 binary values per datapoint). As the confidence intervals suggest, the initial steep increase in correctness after the first trial for the feedback group might be just noise; however, in the second half of the experiment, the feedback group consistently outperforms the no-feedback group.
@@ 3.2.2 Confidence
There is no significant difference in the average confidence between feedback and no-feedback groups (feedback group: 4.19, CI: 4.01–4.35, no-feedback: 4.3, CI: 4.18–4.47), but the dependence of correctness on confidence is different (see Fig 5). At low confidence levels, there was little difference between the groups. However, from approximately level five and above, where the feedback group participants were confident, they indeed showed higher correctness. The greatest difference between the groups appears at confidence level seven, where no feedback group participants were overconfident and made the most errors precisely when they were most certain (even below the chance level of 0.5), while those who could update their assumptions via feedback were justifiably confident, achieving decisively better results.
This difference is also reflected in the model: there is no significant effect in the feedback group, but the more no-feedback group participants were confident in their answer, the worse was their correctness rate (p=0.004; CI -0.23 – -0.045; estimate -0.14).
@@ 3.2.3 Influence of register
Firstly, we look at all the dimensions hollistically, throught Euclidean distance. We utilized the fact that both a human-generated text chunk and an AI-generated text chunk are continuations of the same text, so they should be comparable. It should therefore hold that texts that are stylistically more similar should be harder to distinguish, as this indicates that the AI-generated text chunk is a more credible continuation of the given text. This hypothesis was supported only for no feedback group, where the effect was marginally significant (p = 0.049; estimate: 0.099; CI: 0.00023–0.20). The positive estimate suggests that greater stylistic distance between the texts made it easier for participants to identify which text is the AI-generated continuation. For feedback group, however, the effect was not significant (p = 0.95; estimate: -0.0034; CI: -0.12–0.11).
@@
A similar effect was found for Dim2 difference (p = 0.009; estimate: -0.35; CI: -0.62 – -0.088). That means, when the AI-generated continuation shifted from the spontaneous pole toward the prepared pole, it was more likely to be correctly identified. This finding supports the idea that the no feedback group associate prepared, formal styles — such as those typical of scientific or administrative texts — with AI. When paired with a more spontaneous and subjective human continuation (e.g., featuring dialogue or narrative elements), the contrast likely reinforces the attribution of the more prepared text to AI.
Significant effects were also found in no feedback group for the third and fourth dimensions. When the AI-generated continuation showed a shift from lower to higher cohesion, participants were more likely to correctly identify it as AI-generated (p = 0.19; estimate: 0.098; CI: 0.016–0.18). Similarly, a shift from monothematic to polythematic content was associated with higher accuracy in AI identification (p = 0.16; estimate: 0.12; CI: 0.022–0.22).
@@ 3.2.4 Percieved readability
This finding is also visible in the mixed model: the variable More readable is human has significant results for both groups, but the effect size sil larger for the no feedback group (feedback: p < 0.001; CI 0.68–1.76; estimate 1.22; no feedback: p < 0.001; CI 1.62–2.48; estimate 2.05).
@@ 3.2.5 Attitudes towards AI
In the feedback group, only one statement reached the level of significance: “I prefer to avoid using AI when writing important texts” (p=0.022; CI 0.018–0.23; estimate 0.12). People who agreed with this statement tended to have better correctness rate than people who dissagreed. In the no feedback group, the only statement reaching level of significance was: “AI produces grammatically correct texts.” When people agreed with this statement, they tended to have a better correctness rate than when disagreed (p < 0.001; CI 0.10–0.36; estimate 0.23). This finding is intuitively interpretable: if a participant believes that AI produces grammatically incorrect texts, they may struggle more in this task, as AI-generated texts in our experiment were, in fact, largely grammatically correct.

### v3

@@ Abstract
We used GPT-4o to generate several hundred texts across various genres and text types comparable to Koditex, a multi-register corpus of human-written texts. We then presented randomized text pairs to 254 Czech native speakers who identified which text was human-written and which was AI-generated. Participants were randomly assigned to two conditions: one receiving immediate feedback after each trial, the other receiving no feedback until experiment completion. We recorded accuracy in identification, confidence levels, response times, and judgments about text readability along with demographic data and participants’ engagement with AI technologies prior to the experiment.
@@ 1 Introduction
Participants (n=254) were randomly assigned to two experimental conditions. One group received immediate feedback after each trial, indicating correctness, whereas the second group received no immediate feedback and learned their overall results only upon completing the experiment.
@@ 2.4 Participants
The main sample was primarily recruited from the participant pool managed by the LABELS psycholinguistic laboratory. The pool consists of volunteers interested in participating in linguistic experiments and students who were offered university course credits for their participation. The dataset also includes several participants from the general public who responded to Facebook posts by the experimenters, accounting for approximately 22 participants. The recruitment started 25.02.2025 and ended 14.03.2025. The study was approved by the etical committee of the Charles University beforehand.
The complete sample comprised 291 participants. However, 33 participants were excluded for failing to answer all attention-check questions correctly (the experiment was demanding, and some students participating for course credits showed insufficient attention). An additional 3 participants were excluded for not being native Czech speakers and one for low age (under 18 years old).
The final sample consisted of 254 participants (female: 180, male: 70, other or preferred not to say: 4), the mean age was 24.11, with participants ranging from 19 to 80 years old. The majority of participants had a bachelor’s (99) or master’s (135) degree as their highest level of education, 14 participants were doctoral graduates, and only a few had completed only highschool (3), vocational (2) or other (1) type of education. The majority of participants had an educational background in social sciences and humanities (107), substantially represented were also students and graduates of natural sciences (75). 46 participants specifically indicated that they have philological education, 11 participants studied or graduated in computer science, and 15 participants selected the option ‘Other’.
@@ 3.1 Correctness of the Answer
We observed a substantial and statistically significant difference between feedback and no feedback groups. The no feedback group showed considerably lower performance, with an average correctness rate of 55.4% (95% CI: 52.0%–58.7%), while the feedback group achieved a higher average success rate of 65.1% (95% CI: 62.4%–67.8%). These findings confirm hypotheses 1 and 2: participants were able to distinguish AI-generated texts from human-written ones at above-chance levels, though no feedback group performed only slightly above chance.
@@ 3.2 Mixed effects logistic regression
-2.25in0in     Conf. Interval Conf. Interval Variable     Estimate SE Lower Upper Exp(B) Lower Upper z p (Intercept)     1.07430 0.3503 0.38767 1.76092 2.928 1.474 5.818 3.0666 0.002 Confidence     -0.13657 0.0482 -0.23113 -0.04202 0.872 0.794 0.959 -2.8310 0.005 More readable is human     2.06307 0.2178 1.63617 2.48996 7.870 5.135 12.061 9.4719 <.001 Trial order     -0.01233 0.0113 -0.03447 0.00981 0.988 0.966 1.010 -1.0917 0.275 Reaction time log     -0.08757 0.0832 -0.25057 0.07544 0.916 0.778 1.078 -1.0529 0.292 Dim. Euclidean distance     0.09811 0.0504 -7.66e-4 0.19699 1.103 0.999 1.218 1.9448 0.052 Dim1 difference     -0.30027 0.0850 -0.46679 -0.13376 0.741 0.627 0.875 -3.5344 <.001 Dim2 difference     -0.33332 0.1332 -0.59441 -0.07224 0.717 0.552 0.930 -2.5022 0.012 Dim3 difference     0.09330 0.0415 0.01203 0.17458 1.098 1.012 1.191 2.2500 0.024 Dim4 difference     0.11974 0.0500 0.02172 0.21776 1.127 1.022 1.243 2.3942 0.017 Dim5 difference     -0.08746 0.0543 -0.19392 0.01901 0.916 0.824 1.019 -1.6100 0.107 Dim6 difference     0.00881 0.0346 -0.05902 0.07664 1.009 0.943 1.080 0.2545 0.799 Dim7 difference     0.03550 0.0454 -0.05350 0.12450 1.036 0.948 1.133 0.7817 0.434 Dim8 difference     -0.04195 0.0344 -0.10945 0.02555 0.959 0.896 1.026 -1.2181 0.223 Age     0.01868 0.0139 -0.00859 0.04594 1.019 0.991 1.047 1.3428 0.179 Education     0.08890 0.1292 -0.16431 0.34210 1.093 0.848 1.408 0.6881 0.491 Gender male - female     -0.01249 0.1846 -0.37436 0.34938 0.988 0.688 1.418 -0.0676 0.946 Gender other - female     -0.4545 0.5167 -1.4673 0.5583 0.635 0.231 1.748 -0.88 0.379 Reading disorder     0.71990 0.6945 -0.64126 2.08105 2.054 0.527 8.013 1.0366 0.300 Study field: human - cs     -0.38585 0.3722 -1.11542 0.34372 0.680 0.328 1.410 -1.0366 0.300 Study field: nature - cs     -0.29247 0.3946 -1.06583 0.48089 0.746 0.344 1.618 -0.7412 0.459 Study field: other - cs     -0.36696 0.4746 -1.29717 0.56326 0.693 0.273 1.756 -0.7732 0.439 Study field: philology - cs     -0.51098 0.4184 -1.33099 0.30904 0.600 0.264 1.362 -1.2213 0.222 Frequency of AI usage     0.01728 0.0623 -0.10473 0.13929 1.017 0.901 1.149 0.2777 0.781 AI attitudes: improvement     0.08242 0.0736 -0.06181 0.22666 1.086 0.940 1.254 1.1200 0.263 AI attitudes: risk     -0.02545 0.0580 -0.13910 0.08820 0.975 0.870 1.092 -0.4389 0.661 AI attitudes: quality     -0.01152 0.0670 -0.14286 0.11981 0.989 0.867 1.127 -0.1720 0.863 AI attitudes: creativity     -0.01039 0.0654 -0.13853 0.11774 0.990 0.871 1.125 -0.1590 0.874 AI attitudes: understanding     0.07057 0.0631 -0.05313 0.19427 1.073 0.948 1.214 1.1182 0.263 AI attitudes: grammar     0.23112 0.0645 0.10469 0.35754 1.260 1.110 1.430 3.5830 <.001 AI attitudes: avoidance     -0.07768 0.0581 -0.19155 0.03619 0.925 0.826 1.037 -1.3370 0.181 AI attitudes: dumbing     0.05286 0.0524 -0.04988 0.15559 1.054 0.951 1.168 1.0084 0.313 AI attitudes: fascination     -0.01330 0.0816 -0.17320 0.14661 0.987 0.841 1.158 -0.1630 0.871
@@ 3.2.1 Trial order
This variable is important for our first and second research question. If we are asking whether the group that received feedback differs from the one that did not, we assume that those with feedback were able to learn. Their abilities should therefore improve over time. The model confirmed that they do. The variable Trial order reached level of significance p < 0.001 in feedback group (CI 0.027–0.079). The log odds ratio estimate was at 0.053, that means, the later the trial, the higher probability that the participant answers correctly. In the no feedback group, on the other hand, the trial order did not reach the level of significance (p=0.25, estimate -0.012).
FIG 4 shows that the correctness rises rather chaotically, which is due to the high variability of the data (half of 254 binary values per datapoint). As the confidence intervals suggest, the initial steep increase in correctness after the first trial for the feedback group might be just noise; however, in the second half of the experiment, the feedback group consistently outperforms the no-feedback group.
@@ 3.2.2 Confidence
There is no significant difference in the average confidence between feedback and no-feedback groups (feedback group: 4.2, CI: 4.0–4.4, no-feedback: 4.3, CI: 4.2–4.5), but the dependence of correctness on confidence is different (see Fig 5). At low confidence levels, there was little difference between the groups. However, from approximately level five and above, where the feedback group participants were confident, they indeed showed higher correctness. The greatest difference between the groups appears at confidence level seven, where no feedback group participants were overconfident and made the most errors precisely when they were most certain (even below the chance level of 0.5), while those who could update their assumptions via feedback were justifiably confident, achieving decisively better results.
This difference is also reflected in the model: there is no significant effect in the feedback group, but the more no-feedback group participants were confident in their answer, the worse was their correctness rate (p=0.005; estimate -0.14; CI -0.23 – -0.042).
@@ 3.2.3 Influence of register
Firstly, we look at all the dimensions hollistically, throught Euclidean distance. We utilized the fact that both a human-generated text chunk and an AI-generated text chunk are continuations of the same text, so they should be comparable. It should therefore hold that texts that are stylistically more similar should be harder to distinguish, as this indicates that the AI-generated text chunk is a more credible continuation of the given text. This hypothesis was supported neither for the no feedback group (p = 0.052; estimate: 0.098; CI: 0.000–0.20), neither for the feedback group (p = 0.95; estimate: -0.0034; CI: -0.12–0.11).
@@
A similar effect was found for Dim2 difference (p = 0.012; estimate: -0.33; CI: -0.59 – -0.072). That means, when the AI-generated continuation shifted from the spontaneous pole toward the prepared pole, it was more likely to be correctly identified. This finding supports the idea that the no feedback group associate prepared, formal styles — such as those typical of scientific or administrative texts — with AI. When paired with a more spontaneous and subjective human continuation (e.g., featuring dialogue or narrative elements), the contrast likely reinforces the attribution of the more prepared text to AI.
Significant effects were also found in no feedback group for the third and fourth dimensions. When the AI-generated continuation showed a shift from lower to higher cohesion, participants were more likely to correctly identify it as AI-generated (p = 0.024; estimate: 0.093; CI: 0.012–0.17). Similarly, a shift from monothematic to polythematic content was associated with higher accuracy in AI identification (p = 0.017; estimate: 0.12; CI: 0.022–0.22).
@@ 3.2.4 Percieved readability
This finding is also visible in the mixed model: the variable More readable is human has significant results for both groups, but the effect size is larger for the no feedback group (feedback: p < 0.001; estimate 1.22; CI 0.68–1.76; no feedback: p < 0.001; estimate 2.06; CI 1.64–2.49).
@@ 3.2.5 Attitudes towards AI
In the feedback group, only one statement reached the level of significance: “I prefer to avoid using AI when writing important texts” (p=0.022; estimate 0.12; CI 0.018–0.23). People who agreed with this statement tended to have better correctness rate than people who dissagreed. In the no feedback group, the only statement reaching level of significance was: “AI produces grammatically correct texts.” When people agreed with this statement, they tended to have a better correctness rate than when disagreed (p < 0.001; estimate 0.23; CI 0.10–0.36). This finding is intuitively interpretable: if a participant believes that AI produces grammatically incorrect texts, they may struggle more in this task, as AI-generated texts in our experiment were, in fact, largely grammatically correct.